OpenStudy (anonymous):

if a sales man gets successive gain of 15% and 20% then his actual gain?(ans:38) how??

OpenStudy (anonymous):

let sales be 100 therefore with 15% increase sales it 115 and again 20% increase sales is 138 so net gaun of 38%

OpenStudy (anonymous):

thnkxx..

OpenStudy (anonymous):

another technique to calculate it is 15% + 20% + (15+20)/100 = 35% + 3% = 38%

OpenStudy (anonymous):

why u ignore the 0.5 as it comes fm here(15+20)/100=3.5??i think it may be 15*20/100..

OpenStudy (anonymous):

so sorry 15% + 20% + (15*20)/100 = 35% + 3% = 38%